# Hook-Lengths, Symplectic/Orthogonal Contents and Amdeberhan’s Conjectures

### v1

@@ 1. Introduction
A partition of a nonnegative integer $n$ is a sequence of weakly decreasing positive integers $\lambda=(\lambda_{1},...,\lambda_{l})\vdash n$ satisfying $n=|\lambda|=\sum_{i=1}^{l}\lambda_{i}$, where $l$ is called its length. In general, one can also write $\lambda_{i}=0$ for $i>l$. The conjugation of $\lambda$ is a new partition $\lambda^{t}$ defined by $\lambda^{t}_{i}=\#\{j|\lambda_{j}\geq i\}$ for $i=1,...,\lambda_{1}$. For each pair of positive integers $(i,j)\in\lambda$, i.e., $1\leq j\leq\lambda_{i}$, one can define the following combination numbers related to a given partition $\lambda$. The content is $c(i,j):=j-i$ and the hook-length is $h(i,j):=\lambda_{i}+\lambda^{t}_{j}-i-j+1$. We recommend the books [13, 3] to readers for more interesting properties of partitions. The Amdeberhan’s conjectures concern the so called symplectic contents and orthogonal contents of partitions, which are defined as (see [20])
@@
Stanley [17], inspired by the Nekrasov-Okounkov formula [14], proposed the hook-content formula, which relates a sum over products of hook-lengths and contents to an infinite product formula. The hook-contents are related to the representation theory of general linear groups. Thus it is natural to consider its symplectic and orthogonal counterparts, which should involve symplectic contents and orthogonal contents. These are exactly the Amdeberhan’s conjectures in [1]. Recently, some special cases of Amdeberhan’s conjectures was proved by Amdeberhan-Andrews-Ballantine [2] in terms of studying the set of partitions having non-vanishing symplectic contents. In this paper, we prove the general cases of all conjectures concerning symplectic contents and orthogonal contents proposed by Amdeberhan.
The first result follows
@@
Some special cases of the above two theorems are independently interesting. For example, the Conjecture 6.2 (c) in [1] is the special case $t=0$ of above two formulas and will be used in Corollary (1.4). Next, after establishing a Cauchy-identity type formula for symplectic Schur functions, we prove the following
@@
The Conjecture 6.3 (a) in [1] corresponds to the special case $t_{1}=-t_{2}$ of above Theorem 1.3. See Example 4.4 for more details. As a corollary of Theorem 1.1 and Theorem 1.3, we have
@@
The cases $t=0$ of Theorem 1.1 and Corollary 1.2, and the cases $t_{1}=-t_{2}=0$ and $t_{1}=-t_{2}=1$ of Theorem 1.3 were recently proved by Amdeberhan, Andrews and Ballantine in [2]. Using the method developed in this paper, we can also provide the following new formula, which involves symplectic contents and ordinary contents at the same time,
@@
The rest of this paper is organized as follows. In Section 2, we review the notations of Schur functions, symplectic Schur functions and vertex operators. The latter serves as the main tool in this paper. In Section 3, by studying the partition function consisting of single symplectic Schur functions, we prove Theorem 1.1 and Corollary 1.2. In Section 4, in terms of establishing a Cauchy-identity type formula for symplectic Schur functions, we prove Theorems 1.3, 1.5 and Corollary 1.4.
@@ 2. Preliminaries
In this section, we review the notions of Schur functions, symplectic Schur functions and the vertex operators. We recommend the Macdonald’s book [13] and [5].
@@ 2.1. Schur functions and symplectic Schur functions
The theory of symmetric functions plays important roles in the representation theory. In general, one starts from a set of independent variables $(x_{1},...,x_{n})$ and denote by $\Lambda_{n}(x_{1},...,x_{n})$ the set of symmetric polynomial functions in these variables. It is well-know that the following power sum coordinates
@@
It is more convenient to extend the notion of symmetric functions to functions in infinitely many variables $(\bm{x})=(x_{1},x_{2},...)$. In terms of certain projective limit, one can obtain the ring $\Lambda(\bm{x})$ consisting of symmetric functions in infinitely many variables $(\bm{x})=(x_{1},x_{2},...)$ (for more details, see Chapter I.2 in [13]). Here we concern on some concrete symmetric functions. The $n-$th power sum coordinates $p_{n}(\bm{x})$ are still defined by equation (2.1) but involves infinite many variables $(\bm{x})$. General symmetric functions could also be represented as polynomials in terms of these power sum coordinates. For example, the famous Schur functions can be defined by this way using generating function of elementary Schur functions and Jacobi-Trudi formula. In this paper, we formally define them as follows. The Schur function labeled by a partition $\lambda=(\lambda_{1},...,\lambda_{l})$ is
@@
For defining the symplectic Schur functions, one should consider symmetric functions involving $2n$ variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$. Then, for any partition $\lambda$ satisfying $l(\lambda)\leq n$, the symplectic Schur function labeled by $\lambda$ and $n$ is defined by
@@
It is also known that symplectic Schur function admits the following Jacobi-Trudi formula (see [22, 11], and see [7] for a bijective proof)
@@
where $(x_{1}^{\pm},...,x_{n}^{\pm})=(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$, and $h_{k}(\cdot)$ is the $k-$th complete homogeneous symmetric function in $2n$ variables. Thus, the symplectic Schur function $sp_{\lambda,2n}$ could also be regarded as a symmetric function in variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$. When considering infinitely many variables, we use the following notation
@@
Then the symplectic Schur function $sp_{\lambda}(\tilde{\bm{x}})$ is just defined by the limit of $sp_{\lambda,2n}$. At this case, $sp_{\lambda}(\tilde{\bm{x}})$ could also be represented by a polynomial in those power sum coordinates ${p_{k}(\tilde{\bm{x}})}$.
@@
It is worth mention that, when using the notation $sp_{\lambda,2n}$, there is a hidden condition $l(\lambda)\leq n$. Thus, when considering properties involving symplectic Schur functions labeled by all partitions, it is better to use $sp_{\lambda}$ and regard the corresponding symplectic Schur function as a polynomial in $(\bm{p})$, where $(\bm{p})=(p_{1},p_{2},...)$ could be regarded as formal variables, as well as the $\big{(}p_{k}(\tilde{\bm{x}})\big{)}$.
The importance of Schur functions is that they encode the information of irreducible representations of symmetric groups, as well as the polynomial representations of general linear groups (see [13, 16]). So do symplectic Schur functions. They encodes the information of irreducible polynomial representations of even symplectic groups (see [20]). For example, the following special evaluation of symplectic Schur functions (see [6, 4])
@@
Since both sides of above equation could be regarded as polynomials in $n$ of finite orders and above equation holds for any positive integer $n\geq l(\lambda)$, it should still hold when replacing $n$ by any complex number $t$.
@@ 2.2. Inner product and the vacuum expectation value
There is a standard inner product on the ring $\mathbb{C}[p_{1},p_{2},...]$ by letting $(s_{\lambda}(\bm{p}))_{\lambda\in\mathcal{P}}$ be a standard normalized basis.
Below, we introduce the notation of vacuum expectation value, which is more convenient (see, for example, [5]). For avoiding the misuse of symbols, we use the $(\bm{q})=(q_{k})$ coordinates and vectors appeared in vacuum expectation value are regarded as elements in $\mathbb{C}[\bm{q}]$. First, denote by $|\lambda\rangle$ the Schur function $s_{\lambda}(\bm{q})$ when regarding it as a vector in the vector space $\mathbb{C}[\bm{q}]$, and $\langle\lambda|$ is its dual. The vacuum $|0\rangle$ denotes the vector corresponding to the Schur function labeled by empty partition, which is the constant function 1, and $\langle 0|$ is its dual. Then the vacuum expectation value is of the following form
@@ 2.3. Vertex operator
In this subsection, we review the notion of vertex operator. It is an important tool to study the Schur functions and related symmetric functions. We mainly follow the notation in [5].
Vertex operators are generating functions of differential operators on the ring $\mathbb{C}[\bm{q}]$. First, denote by the Heisenberg operators
@@
where $q_{-n}\cdot$ means the operator multiplying by $q_{-n}$, then the vertex operators are directly defined by
@@
When regarding $(\bm{p})$ as the power sum coordinates $\big{(}\bm{p}(\bm{y})\big{)}=\big{(}p_{k}(\bm{y})\big{)}$ in symmetric variables $(\bm{y})=(y_{1},y_{2},...)$, one has
@@
The special cases of above equations give
@@
which could also be proved from their definitions obviously.
3. Partition function consisting of single symplectic Schur functions
In this section, by studying the partition function consisting of single symplectic Schur functions, we prove Theorem 1.1 and Corollary 1.2.
3.1. Partition function of single symplectic Schur functions
@@
Both sides of above equation could be regarded as elements in the ring $\mathbb{C}[\![q]\!][\bm{p}]$.
Above theorem has the following equivalent formalism expressed in terms of symmetric variables $(\bm{y})=(y_{i})$.
@@
Write $p_{k}$ as $p_{k}(\bm{y})$, the $k-$power sum of variables $(\bm{y})$. Then Theorem 3.1 is equivalent to the following
@@
Then the right hand side of equation (3.2) is a direct rewriting of equation (3.1) in terms of symmetric variables $(\bm{y})=(y_{i})$. $\Box$
@@
In this subsection, we prove Theorem 3.1 by showing equation (3.2).
First, we need to introduce the following operator
@@
The operators $\Gamma_{-}(z)$ and $\exp(G)$ commute to each others. For $\Gamma_{+}(z)$, we have the following commutation relations
@@
Proof: The first result that $\Gamma_{-}(z)$ and $\exp\Big{(}\sum_{n=1}\frac{1}{n}(\frac{1}{2}\alpha_{-n}^{2}-\frac{1}{2}\alpha_{-2n})\Big{)}$ commute to each others follows from the fact that both of them only consist of $\alpha_{-n}$ for $n>0$. The equation (3.5) follows from the following two commutation relations
@@
Above two equations can be proved by using Baker-Campbell-Hausdorff formula and for the second equation, we also need the following
@@
which is proved by Zassenhaus formula. The equation (3.6) can be proved by multiplying $\Gamma_{+}(z)^{-1}$ to both sides of equation (3.5). $\Box$
@@
For vertex operator $\Gamma_{+}\big{(}\bm{p}(\bm{y})\big{)}$ defined in equation (2.7), we have the following commutation relations
@@
Proof: Just recalling that $\Gamma_{+}\big{(}\bm{p}(\bm{y})\big{)}=\prod_{j=i}^{\infty}\Gamma(y_{i})$, and using Lemma 3.3, we have
@@
Then, this lemma follows from the commutation relation (2.5). $\Box$
@@
Denote by $F^{*}$ its dual. Then we have
@@
Operator $\exp(F^{*})$ commutes to $\Gamma_{+}(\bm{p}(\bm{y}))$. For $\Gamma_{-}(\bm{p}(\bm{y}))$, we have
@@
Proof: It was proved by the author as Lemma 5.2 in [24] that
@@
By using the dual of above equation, one can prove
@@
Then equation (3.10) follows from the above equation and the commutation relation (2.5). For proving equation (3.11), one only needs to multiply $\Gamma\big{(}\bm{p}(\bm{y})\big{)}^{-1}$ to both sides of equation (3.10). $\Box$
With above commutation relations at hands, we are prepared to prove Theorem 3.1.
Proof of Theorem 3.1: Denote by $\mathcal{P}^{\prime}$ the set of partitions whose Frobenius notations are of the form $(m_{1},...,m_{r}|m_{1}+1,...,m_{r}+1)$. For convenience, $\emptyset$ is in $\mathcal{P}^{\prime}$. The following result was proved by Sundaram in her thesis (see Theorem 12.10 in [18])
@@
which is motivated by a classical result representing Schur functions in terms of symplectic Schur functions obtained by Littlewood (see, for example, [12]). Sundaram provided equation (3.12) when regarding both sides of above equation as formal functions in variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$ for $l(\lambda)\leq n$. Here, equation (3.12) holds as an equality in the ring $\mathbb{C}[\bm{p}]$.
Then by the definition (2.2) of skew Schur functions, we have
@@
where in the second equal sign, we have used the formula (2.8), which expresses skew Schur function in terms of vacuum expectation value involving vertex operator.
The following equation is useful (see I.5.Example 9 (a) in [13]), which is the motivation to introduce the operator $G$,
@@
In terms of the power sum coordinates $\big{(}\bm{p}(\bm{y})\big{)}$, one has
@@
Then, by the definition equation (3.4) of the operator $G$, above two equations can be rewritten as
@@
For the operator $F$, we have (see I.5.Example 4 in [13], see also equation (5.4) in [24]),
@@
From now on, we replace the variables $(\bm{p})$ by symmetric functions $\big{(}\bm{p}(\bm{y})\big{)}$ for convenience. One can first use equation (3.10) to change the orders of the operators $\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}$ and $\exp(F^{*})$. The result is that, $\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}$ becomes $\Gamma_{+}\big{(}\bm{p}(q^{2}\bm{y})\big{)}$ since $\langle 0|\Gamma_{-}(\cdot)=\langle 0|$. Then one can use equation (3.7) to deal with the commutation relation between operators $\Gamma_{+}\big{(}\bm{p}(q^{2}\bm{y})\big{)}$ and $\exp(G)$. More precisely, above equation (3.16) reduces to that, the left hand side of it is equal to
@@
where we have used the fact $\Gamma_{+}(\cdot)|0\rangle=|0\rangle$. Similarly, One can use the equations (3.11) and (3.8) to change the orders of the operators $\Gamma_{\pm}(\cdot)^{-1}$, $\exp(F^{*})$ and $\exp(G)$. The result is that, the left hand side of equation (3.16) is equal to
@@
One can compare the above equation and equation (3.16). Thus, for an arbitrary positive integer $M$, one can repeat the above process $M$ times to obtain the following
@@
Since both sides of above equation are regarded as elements in the ring $\Lambda(\bm{y})[\![q]\!]$, we have, as operator on the ring $\Lambda(\bm{y})[\![q]\!]$,
@@
where Id means the identity operator. Thus, this theorem is proved by taking the limit $M\rightarrow+\infty$ of equation (3.17) and the following combinatorial result
@@
where the first equal sign follows from equations (3.14) and (3.15) and the second equal sign follows from the fact that the size of the partition $(m_{1},...,m_{r}|m_{1}+1,...,m_{r}+1)\in\mathcal{P}^{\prime}$ is equal to $2\sum_{i=1}^{r}(m_{i}+1)$. $\Box$

### v2

@@ 1. Introduction
A partition of a nonnegative integer $n$ is a sequence of weakly decreasing positive integers $\lambda=(\lambda_{1},...,\lambda_{l})\vdash n$ satisfying $n=|\lambda|=\sum_{i=1}^{l}\lambda_{i}$, where $l$ is called its length. In general, one can also write $\lambda_{i}=0$ for $i>l$. The conjugation of $\lambda$ is a new partition $\lambda^{t}$ defined by $\lambda^{t}_{i}=\#\{j|\lambda_{j}\geq i\}$ for $i=1,...,\lambda_{1}$. For each pair of positive integers $(i,j)\in\lambda$, i.e., $1\leq j\leq\lambda_{i}$, one can define the following combination numbers related to the given partition $\lambda$. The content is $c(i,j):=j-i$ and the hook-length is $h(i,j):=\lambda_{i}+\lambda^{t}_{j}-i-j+1$. We recommend the books [3, 13] to readers for more interesting properties of partitions. The Amdeberhan’s conjectures concern the so called symplectic contents and orthogonal contents of partitions, which are defined as (see [20])
@@
Stanley [17], inspired by the Nekrasov-Okounkov formula [14], proposed the hook-content formula, which relates a sum over products of hook-lengths and contents to an infinite product formula. The hook-contents are related to the representation theory of general linear groups. Thus it is natural to consider its symplectic and orthogonal counterparts, which should involve symplectic contents and orthogonal contents. These are exactly the Amdeberhan’s conjectures in [1].
Our first result is stated as follows
@@
Some special cases of the above two results are independently interesting. For example, Conjecture 6.2 (c) in [1] is the special case $t=0$ of the above two formulas and will be used in Corollary 1.4. Next, after establishing a Cauchy-identity type formula for symplectic Schur functions, we prove the following
@@
Conjecture 6.3 (a) in [1] corresponds to the special case $t_{1}=-t_{2}$ of Theorem 1.3. See Example 4.4 for more details. As a consequence of Theorem 1.1 and Theorem 1.3, we have
@@
The cases $t=0$ of Theorem 1.1 and Corollary 1.2, and the cases $t_{1}=-t_{2}=0$ and $t_{1}=-t_{2}=1$ of Theorem 1.3 were recently proved by Amdeberhan, Andrews and Ballantine in [2]. Using the method developed in this paper, we can also provide the following new formula, which involves symplectic contents and ordinary contents at the same time.
@@
The rest of this paper is organized as follows. In Section 2, we review the concepts of Schur functions, symplectic Schur functions and vertex operators. The latter serves as the main tool in this paper. In Section 3, by studying a partition function consisting of symplectic Schur functions, we prove Theorem 1.1 and Corollary 1.2. In Section 4, after establishing a Cauchy-identity type formula for symplectic Schur functions, we prove Theorems 1.3, 1.5 and Corollary 1.4.
@@ 2. Preliminaries
In this section, we review the concepts of Schur functions, symplectic Schur functions and the vertex operators. We recommend the Macdonald’s book [13] and [5].
@@ 2.1. Schur functions and symplectic Schur functions
The theory of symmetric functions plays important roles in the representation theory. In general, one starts from a set of independent variables $(x_{1},...,x_{n})$ and denote by $\Lambda_{n}(x_{1},...,x_{n})$ the set of symmetric polynomial functions in these variables. It is well-known that the following power sum coordinates
@@
It is more convenient to extend the notion of symmetric functions to functions in infinitely many variables $(\bm{x})=(x_{1},x_{2},...)$. In terms of certain projective limit, one can obtain the ring $\Lambda(\bm{x})$ consisting of symmetric functions in infinitely many variables $(\bm{x})=(x_{1},x_{2},...)$ (for more details, see Chapter I.2 in [13]). Here we are focusing on some concrete symmetric functions. The $n$-th power sum coordinates $p_{n}(\bm{x})$ are still defined by equation (2.1) but they now involve infinitely many variables $(\bm{x})$. General symmetric functions could also be represented as polynomials in terms of these power sum coordinates. For example, the famous Schur functions can be defined in this manner by using generating function of elementary Schur functions and Jacobi-Trudi formula. In this paper, we formally define them as follows. The Schur function labeled by a partition $\lambda=(\lambda_{1},...,\lambda_{l})$ is
@@
To introduce the symplectic Schur functions, one should consider symmetric functions involving $2n$ variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$. Then, for any partition $\lambda$ satisfying $l(\lambda)\leq n$, the symplectic Schur function labeled by $\lambda$ and $n$ is defined by
@@
It is also known that the symplectic Schur function admits the following Jacobi-Trudi formula (see [11, 22], and see [7] for a bijective proof)
@@
where $(x_{1}^{\pm},...,x_{n}^{\pm})=(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$, and $h_{k}(\cdot)$ is the $k$-th complete homogeneous symmetric function in $2n$ variables. Thus, the symplectic Schur function $sp_{\lambda,2n}$ could also be regarded as a symmetric function in the variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$. In dealing with infinitely many variables, we adopt the following notation
@@
Then the symplectic Schur function $sp_{\lambda}(\tilde{\bm{x}})$ is just formally defined by the limit $n\rightarrow\infty$ of $sp_{\lambda,2n}$. In this case, $sp_{\lambda}(\tilde{\bm{x}})$ could also be represented by a polynomial in those power sum coordinates ${p_{k}(\tilde{\bm{x}})}$ and we write it as $sp_{\lambda}(\bm{p})$, where $(\bm{p})=(p_{1},p_{2},...)$, if there is no ambiguity.
@@
It is worth mentioning that, when using the notation $sp_{\lambda,2n}$, there is a hidden condition $l(\lambda)\leq n$. Thus, when considering properties involving symplectic Schur functions labeled by all partitions, it is better to use $sp_{\lambda}$ and regard the corresponding symplectic Schur function as a polynomial in $(\bm{p})$.
The importance of Schur functions is that they encode information on irreducible representations of symmetric groups, as well as the polynomial representations of general linear groups (see [13, 16]). So do symplectic Schur functions. They encodes information on the irreducible polynomial representations of even symplectic groups (see [20]). For example, the following special evaluation of the symplectic Schur functions (see [4, 6])
@@
Since both sides of the above equation could be treated as polynomials in $n$ of finite orders, and since the above equation holds for any positive integer $n\geq l(\lambda)$, it should still be valid when replacing $n$ by any complex number $t$.
@@ 2.2. Inner product and the vacuum expectation value
There is a standard inner product on the ring $\mathbb{C}[p_{1},p_{2},...]$ by letting $(s_{\lambda}(\bm{p}))_{\lambda\in\mathcal{P}}$ be a standard normalized basis (see [13] for more details).
Below, we introduce the notation of vacuum expectation value, which is more convenient for our purposes (see, for example, [5]). To avoid the misuse of symbols, we employ the $(\bm{q})=(q_{k})$ coordinates and vectors that appear in vacuum expectation value are viewed as elements in $\mathbb{C}[\bm{q}]$. First, denote by $|\lambda\rangle$ the Schur function $s_{\lambda}(\bm{q})$ when regarding it as a vector in the vector space $\mathbb{C}[\bm{q}]$, and $\langle\lambda|$ is its dual. The vacuum $|0\rangle$ denotes the vector corresponding to the Schur function labeled by the empty partition, which is the constant function 1, and $\langle 0|$ is its dual. Then the vacuum expectation value is of the following form
@@ 2.3. Vertex operator
In this subsection, we review the concept of vertex operators. It is an important tool in studying the Schur functions and related symmetric functions. We mainly follow the notations in [5].
Vertex operators are generating functions of differential operators on the ring $\mathbb{C}[\bm{q}]$. First, introduce the Heisenberg operators
@@
where $q_{-n}\cdot$ means the operation of multiplying by $q_{-n}$. Then, the vertex operators can be defined directly by
@@
When regarding $(\bm{p})$ as the power sum coordinates $\big{(}\bm{p}(\bm{y})\big{)}=\big{(}p_{k}(\bm{y})\big{)}$, in the symmetric variables $(\bm{y})=(y_{1},y_{2},...)$, one has
@@
The special cases of the above equations give
@@
which could easily be proven from their definitions.
3. A partition function consisting of symplectic Schur functions
In this section, by studying a partition function consisting of symplectic Schur functions, we prove Theorem 1.1 and Corollary 1.2.
3.1. A partition function of symplectic Schur functions
@@
Both sides of the above equation could be regarded as elements in the ring $\mathbb{C}[\![q]\!][\bm{p}]$.
Theorem 3.1 has the following equivalent formalism expressed in terms of the symmetric variables $(\bm{y})=(y_{i})$.
@@
Write $p_{k}$ as $p_{k}(\bm{y})$, the $k$-power sum of variables $(\bm{y})$. Then, Theorem 3.1 is equivalent to the following
@@
Then the right hand side of equation (3.2) is a direct rewriting of equation (3.1) in terms of the symmetric variables $(\bm{y})=(y_{i})$. $\Box$
@@
In this subsection, we prove Theorem 3.1 by showing equation (3.2). First, we need to introduce the following operator
@@
The operators $\Gamma_{-}(z)$ and $\exp(G)$ commute with each other. For $\Gamma_{+}(z)$, we have the following commutation relations
@@
Proof: The first result that $\Gamma_{-}(z)$ and $\exp\Big{(}\sum_{n=1}\frac{1}{n}(\frac{1}{2}\alpha_{-n}^{2}-\frac{1}{2}\alpha_{-2n})\Big{)}$ commute with each other follows from the fact that both of them consist only of $\alpha_{-n}$ for $n>0$. Equation (3.5) is a consequence of the following two commutation relations
@@
In return, these two equations can be proved by using the Baker-Campbell-Hausdorff formula while the second equation additionally demands the following
@@
which is provable by the Zassenhaus formula. Equation (3.6) can be shown by multiplying both sides of equation (3.5) by $\Gamma_{+}(z)^{-1}$. $\Box$
@@
For vertex operator $\Gamma_{+}\big{(}\bm{p}(\bm{y})\big{)}$, defined in equation (2.7), we have the following commutation relations
@@
Proof: Just recalling that $\Gamma_{+}\big{(}\bm{p}(\bm{y})\big{)}=\prod_{j=i}^{\infty}\Gamma(y_{i})$ and using Lemma 3.3, we have
@@
Then, the assertions follow from the commutation relation (2.5). $\Box$
@@
Denote its dual by $F^{*}$. Then we have
@@
The operators $\Gamma_{+}(\bm{p}(\bm{y}))$ and $\exp(F^{*})$ commute with each other. For $\Gamma_{-}(\bm{p}(\bm{y}))$, we have
@@
Proof: The author proved (see Lemma 5.2 in [24]) that
@@
By using the dual of the above equation, one can show
@@
Thus, equation (3.10) follows from the above equation and the commutation relation (2.5). To prove the relation in (3.11), one only needs to multiply both sides of equation (3.10) by $\Gamma\big{(}\bm{p}(\bm{y})\big{)}^{-1}$. $\Box$
With above commutation relations at hand, we are prepared to prove Theorem 3.1.
Proof of Theorem 3.1: Denote by $\mathcal{P}^{\prime}$ the set of partitions whose Frobenius notations are of the form $(m_{1},...,m_{r}|m_{1}+1,...,m_{r}+1)$ for some decreasing non-negative integers $m_{1},...,m_{r}$. For convenience, the empty partition $\emptyset$ is considered in $\mathcal{P}^{\prime}$. The following result was proved by Sundaram in her thesis (see Theorem 12.10 in [18])
@@
which is motivated by a classical result representing Schur functions in terms of symplectic Schur functions obtained by Littlewood (see, for example, [12]). Sundaram provided the equation (3.12) by regarding both sides of the equation as formal functions in variables $(x_{1},x_{1}^{-1},...,x_{n},x_{n}^{-1})$ when $l(\lambda)\leq n$. Here, equation (3.12) is an equality in the ring $\mathbb{C}[\bm{p}]$.
Then from the definition (2.2) of skew Schur functions, we have
@@
where the second equal sign is an application of formula (2.8), which expresses skew Schur function in terms of the vacuum expectation value involving vertex operators.
The following equation is useful (see I.5.Example 9 (a) in [13]), which is in fact our motivation to introduce the operator $G$,
@@
When expressed in terms of the power sum coordinates $\big{(}\bm{p}(\bm{y})\big{)}$, one has
@@
Based on the definition (3.4) of the operator $G$ and the homogeneous operator $L_{0}$, the above two equations can be rewritten as
@@
For the operator $F$, we have the following similar formulas (see I.5.Example 4 in [13], and see also equation (5.4) in [24]),
@@
From now on, and for convenience, we replace the variables $(\bm{p})$ by symmetric functions $\big{(}\bm{p}(\bm{y})\big{)}$. One can first use equation (3.10) to change the orders of the operators $\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}$ and $\exp(F^{*})$. The result is that, $\Gamma_{-}\big{(}\bm{p}(\bm{y})\big{)}$ becomes $\Gamma_{+}\big{(}\bm{p}(q^{2}\bm{y})\big{)}$ since $\langle 0|\Gamma_{-}(\cdot)=\langle 0|$. Then one can use equation (3.7) to deal with the commutation relation between operators $\Gamma_{+}\big{(}\bm{p}(q^{2}\bm{y})\big{)}$ and $\exp(G)$. More precisely, the left hand side of equation (3.16) is equal to
@@
where we have used the fact that $\Gamma_{+}(\cdot)|0\rangle=|0\rangle$. Similarly, one can use equations (3.11) and (3.8) to change the orders of the operators $\Gamma_{\pm}(\cdot)^{-1}$, $\exp(F^{*})$ and $\exp(G)$. The result is that, the left hand side of equation (3.16) is equal to
@@
This time, we compare the above equation and equation (3.16). One can notice that, for an arbitrary positive integer $M$, by repeating the preceding process $M$ times, the following equation can be obtained
@@
Since both sides of the above equation are regarded as elements in the ring $\Lambda(\bm{y})[\![q]\!]$, we have, as operators on the ring $\Lambda(\bm{y})[\![q]\!]$,
@@
where Id denotes the identity operator. Thus, the theorem is proved by taking the limit $M\rightarrow+\infty$ in equation (3.17) and by applying the following combinatorial result
@@
where the first equality follows from equations (3.14) and (3.15), while the second equality follows from the fact that the size of the partition $(m_{1},...,m_{r}|m_{1}+1,...,m_{r}+1)\in\mathcal{P}^{\prime}$ is equal to $2\sum_{i=1}^{r}(m_{i}+1)$. $\Box$